OpenStudy (anonymous):

The length of the base of a square pyramid is 9 inches. The slant height is 12 inches. What is the surface area of the pyramid? A)189 in2 B)225 in2 C)297 in2 D)360 in2

OpenStudy (igreen):

Formula: \(SA_{square~pyramid} = b^2 + 2 \times b \times s\) Where 'b' is the length of the base, and 's' is the slant height. \(SA_{square~pyramid} = 9^2 + 2 \times 9 \times 12\) Simplify

OpenStudy (anonymous):

thanks!

OpenStudy (anonymous):

i get 297